What is the common difference in this sequence?

5
21
37
17
21
29

Answer by Edwin McCravy(20077)

What is the common difference in this sequence?

5
21
37
17
21
29
There is no common difference in them.  They have to be
spaced evenly apart to have a common difference.

21-5 = 16
37-21 = 16

So the sequence of the first three only

5 
21
37 

have a common difference of 16.

But the next number would have to be 53 so that

53-37 = 16

But it isn't.

I don't know why your teacher would expect you to find
a common difference when there is none.

Edwin